OpenStudy (anonymous):

if the number in bacteria in a culture doubles every three hours the bacteria growth can be represented by the formal N(t)=n0(2)t/3 where t is the time in hours and n0 is the number present at the start of the experiment. if 400 bacteria are counted at the start of the experiment how many would be present after 24 hours

OpenStudy (anonymous):

\[N(24)=400\times 2^{\frac{24}{3}}\] will do it or \[N(24)=400\times 2^8\] or even \[N(24)=400\times 256\]